# Bayesian mitigation of measurement errors in multi-qubit experiments

### v1

@@ Abstract
We introduce an implementation of Bayesian measurement error mitigation tailored for multiqubit experiments on near-term quantum devices. Our approach leverages complete information from the readout signal, which is available before any binary state assignment of the qubits. We provide a detailed algorithm workflow, from the calibration of detector response functions to the post-processing of measurement outcomes, offering a computationally efficient solution suitable for the output size typical of current quantum computing devices and quantum algorithms. We benchmark our protocol on actual quantum computers with superconducting qubits, where the readout signal encodes the measurement information in the IQ clouds before qubit state assignment. Finally, we compare the performance of our algorithm against other measurement error mitigation methods, such as iterative Bayesian unfolding and noise matrix inversion.
@@ I Introduction
Despite the rapid advancements in quantum computing technologies, practical applications of noisy intermediate-scale quantum (NISQ) processing units (QPUs) are still out of reach. While QPU manufacturers keep improving quality metrics [1, 2, 3], be it decoherence times, gate and readout fidelities, a lot of effort is going into developing software tools to shrink the gap between acceptable noise levels and utility-scale algorithms [4, 5]. Quantum error mitigation (QEM) embodies such a goal, even though a comprehensive and universally accepted definition for the field is lacking [6]. In general, QEM encompasses all those algorithmic schemes that reduce noise and errors in the expectation value of some figure of merit, such as the average of observables or quantum state probabilities. This is typically achieved by post-processing the outputs from an ensemble of circuit runs, either the same circuit or a family of properly designed ones [7]. This statistical approach starkly contrasts with quantum error correction, which aims to reduce noise after each shot. Among the most advanced techniques, zero-noise extrapolation [8] with probabilistic error amplification and probabilistic error cancellation [9] have recently been found successful in the largest available QPUs [10, 11, 12]. Some specialised techniques aim to target specific sources of errors, such as measurement error mitigation strategies. These strategies are designed to improve the readout step of quantum algorithms execution by modeling the measurement error, which is a result of the faulty physical measurement process. Most readout mitigation techniques rely on detector or measurement tomography [13, 14, 15, 16] to obtain the assignment, noise, or confusion matrix. The readout error is then addressed in various ways. One straightforward strategy involves manipulating the outcome count statistics by multiplying the inverse of the confusion matrix with the measured quantum state probability vector [17]. However, this approach can yield non-physical values, such as negative probabilities, an issue that can be accounted for by incorporating a convex optimizer into the mitigation pipeline [18]. Alternatively, some proposals have developed machine learning-based techniques to perform the assignment process [19, 20, 21, 22].
@@
Figure 1: Pictorial sketch of the heuristic algorithm to perform Bayesian measurement error mitigation in a multiqubit experiment. In the first step (I), we collect the noisy bitstrings counts and reduce the $2^{\mathrm{N_{q}}}$ outcome subspace to mitigate to the $M$ measured noisy counts for which we have at least one measurement. In step (II), we select two populations to use as variables while keeping all others constant, and apply Bayesian measurement error mitigation to obtain two new mitigated values. The Bayesian update is performed for all possible pairs. Step (II) is repeated until a specified convergence criterion or exit condition is met, resulting in the full set of mitigated values (III).
@@
In a recent work [23], we have introduced an efficient and accurate readout measurement scheme for single and multiqubit states based on Bayesian inference, inspired by successful application of Bayesian inference in quantum metrology and sensing [24, 25] or parameter estimation for quantum circuits and states [26, 27, 14, 28, 29, 30, 31, 32]. Our method, bypassing the need to invert the noise matrix, leverages the knowledge of the detector faulty response and Bayesian inference to obtain a probability distribution for each qubit state. Moreover, such an approach is not limited to use binary outcomes from the detector, but it can exploit the detector unprocessed analog readout data, before any binary assignment is made. The idea of using analog data has found traction also in quantum error correction, where it has been shown to lead to improved decoders and logical qubits error rates [33, 34]. Still, all measurement error mitigation schemes suffer some form of scalability issue when moving to many-qubit systems, as intuitively the noise matrix size scales exponentially with the number of qubits. Often this issue is tackled by not using the full assignment matrix, or its inverse, and instead working in a subspace defined by the noisy outcomes [35]. In this work, we aim to tackle thoroughly the scaling problem when aiming to mitigate measurement errors with our formerly introduced Bayesian readout and further show how the use of the readout analog data leads to increased performances. Beyond showing a solid proof of principle of the method, we make a full scaling analysis of the algorithm and provide an estimate of the classical computation time needed to post-process the measurement results obtained from currently available superconducting QPUs.
@@
The paper is organised as follows. In section II, we introduce all the necessary theoretical concepts to model measurement errors in quantum computers and how a Bayesian approach can be used to infer the correct outcomes from the noisy ones, provided the knowledge of the detector noise matrix. Furthermore, we introduce a novel algorithm to implement the Bayesian measurement error mitigation scheme when dealing with multi-qubit measurements. We provide a complete step-by-step procedure which is efficient and feasible for current quantum computing devices, and show how to adapt it for different type of available measurement data, being it binary outcome, i. e. ”0”s or ”1”s, or raw analog readout detector signal, e.g. IQ-clouds for superconducting QPUs.
@@
In section III, we benchmark our implementation of Bayesian measurement error mitigation running quantum algorithms on an actual quantum computer. We show how the algorithm can be efficiently applied to improve readout accuracy for multi-qubit experiments and how the performances of the mitigation scheme scale with the number of qubit and the number of measurements.
In section IV, we focus on the Bayesian measurement error mitigation scheme applied to the detector analog data, showing how using the IQ-readout signal, instead of a faulty binary outcome for each qubit, can improve dramatically the performances of the mitigation algorithm in a multi-qubit scenario. In order to highlight the merit and qualities of our implementation of Bayesian measurement error mitigation, in section V, we compare its performances against two other measurement error mitigation techniques, i.e. the iterative Bayesian unfolding and the noise matrix inversion with positivity constraints. In section VI, we discuss the theoretical computational cost of our implementation of the algorithm for multi-qubit Bayesian measurement error mitigation and give, as an example, the actual classical CPU time spent to perform the mitigation routine for some of the results shown the previous sections. Throughout this work, we run our quantum algorithms on two superconducting qubit devices, the superconducting qubits QPUs VTTQ20 (Leena from IQM) and IBM127 (Nazca). Finally, we present the conclusions and outline some possible future research directions and applications in Section VII.
@@ II Theoretical background
Let’s start with a mathematical description of a noisy measurement process for a single qubit. Let $s\in\{0,1\}$ be the measurement outcome observed in a single experiment. In the absence of measurement errors, the probability to register either one is fully determined by the quantum state of the system. Formally, a conventional single qubit measurement is modelled with a Positive-Operator Valued Measure (POVM) with two effects $\{M_{0}=|0\rangle\langle 0|,M_{1}=|1\rangle\langle 1|\}$. Consequently, the probability of observing a measurement outcome $s$, when the quantum system is in a state $\hat{\rho}$, can be calculated as $\mathrm{Pr}(s)=\operatorname{Tr}(\hat{\rho}M_{s})\equiv\rho_{s}$.
@@
However, in real a setup, the readout step is noisy and errors emerge in the process of collecting the signal from the detector and in the assignment of the correct qubit state. In practice, we can model the experimental noisy measurement trough a set of noisy effects $M^{\mathrm{noisy}}_{i}=\sum_{j}\langle i|\Lambda|j\rangle M_{j}=\sum_{j}\Lambda_{ij}M_{j}$, where $\Lambda$ is the noise matrix and $\Lambda_{ij}$ is the probability of observing/assigning outcome $i$ when the true outcome is $j$. Therefore, after repeated measurements, we have direct access to the noisy probabilities
$\rho^{\mathrm{noisy}}_{s}=\sum_{ij}\Lambda_{ij}\rho_{j}\delta(i-s),$ (1)
@@
which are connected to the true qubit state probabilities trough the noise matrix $\Lambda$. Thus, the goal of measurement error mitigation is to reconstruct the true qubit state probabilities after measuring $\rho^{\mathrm{noisy}}_{s}$. Assuming that we have access to the noise matrix $\Lambda$, from Eq. (1), it is straightforward to attempt to invert the noise matrix and use it to reconstruct the correct qubit state probability vector as $\rho=\Lambda^{-1}\rho^{\mathrm{noise}}$. However, this simple measurement error mitigation strategy might result in a probability vector containing negative values due to statistical errors. This issue is often accounted for by employing a convex optimizer which project the corrected vector to a physical one, even though the issue and the extra step required become more concerning and demanding when dealing with multi-qubit states.
@@
To avoid these issues, we propose here a strategy to bypass the need to invert the noise matrix by noting how the measured outcome $\rho^{\mathrm{noisy}}_{s}$, as introduced in Eq. (1), can be read as the conditional probability of measuring $s$ when the correct qubit state probabilities are $\rho_{0}$ and $\rho_{1}$, i.e. $\rho^{\mathrm{noisy}}_{s}=\mathrm{Pr}(s|\rho_{0},\rho_{1})$.
@@
Thus, we can reconstruct a probability distribution for the correct probabilities $\rho_{0}$ and $\rho_{1}$ trough Bayes’ theorem as $\mathrm{Pr}(\rho_{0},\rho_{1}|s)\propto\left[\Lambda_{s0}\rho_{0}+\Lambda_{s1}\rho_{1}\right]\mathrm{Pr}(\rho_{0},\rho_{1})$. It is convenient to follow an iterative approach and consider individually the $\mathrm{N}_{\mathrm{shots}}$ values collected after each measurement. In this way, we can update the conditional probability for the quantum state probabilities $\rho_{0}$ and $\rho_{1}$ shot after shot as
$\mathrm{Pr}(\rho_{0},\rho_{1}|s^{(n)})\propto\left[\Lambda_{s0}\rho_{0}+\Lambda_{s1}\rho_{1}\right]\mathrm{Pr}(\rho_{0},\rho_{1}|s^{(n-1)}),$ (2)
@@
where $s^{(n)}\in\{0,1\}$ is the bit registered in the n-th experiment, and $\mathrm{Pr}(\rho_{0},\rho_{1}|s^{(n-1)})$ is the prior probability distribution which corresponds to the posterior distribution calculated at the previous step. The idea is to update the prior probability distribution after each iteration step with the newly calculated posterior probability distribution. When assuming no prior knowledge about the system state, we start the iteration with a uniform prior probability distribution $\mathrm{Pr}(\rho_{0},\rho_{1}|s^{(0)})=\mathrm{const}$.
[…]
@@
With these ingredients, we can extend the Bayesian measurement error mitigation scheme to the multi-qubit case with relative ease. For a system of $\mathrm{N}_{q}$ qubits, the measurement outcome obtained from a single experiment is a $\mathrm{N}_{q}$ digits string, $\mathbf{s}\in\{0,1\}^{\mathrm{N}_{q}}$. Equivalent to Eq. (1), the conditional probability of observing the multi-qubit string $\mathbf{s}$ is related to the correct outcome trough a multi-qubit noise matrix as
$\mathrm{Pr}(\mathbf{s}|\bm{\rho}\equiv\rho_{0},...,\rho_{2^{N_{q}}-1})=\sum_{\mathbf{ij}}\Lambda_{\mathbf{ij}}\rho_{\mathbf{j}}\delta(\mathbf{i}-\mathbf{s}).$ (3)
@@
$\mathrm{Pr}(\bm{\rho}|\mathbf{s}^{(n)})\propto\sum_{\mathbf{ij}}\Lambda_{\mathbf{ij}}\rho_{\mathbf{j}}\delta(\mathbf{i}-\mathbf{s}^{(n)})\mathrm{Pr}(\bm{\rho}|\mathbf{s}^{(n-1)}),$ (4)
@@
For example, for a system with ${\mathrm{N_{q}}}$ qubits, Eq. (4) requires handling a multi-parameter probability distribution $\mathrm{Pr}(\bm{\rho}|\mathbf{s}^{(n)})$ at each iteration step of the Bayesian update. This distribution involves $2^{\mathrm{N_{q}}}$ variables $\rho_{\mathbf{i}}$, each sampled within the interval $[0,1]$. If each variable is sampled with $n_{p}$ points, the total probability distribution would necessitate a mesh-grid with $n_{p}^{2^{\mathrm{N_{q}}}}$ points. Even though the normalization condition reduces the effective number of independent coordinates the computational cost remains high and out of reach.
Additional complications arise from the noise matrix resulting of the multi-qubit measurement process. Ideally, the reconstruction of the multi-qubit noise matrix appearing in Eq. (4) would require a full detector tomography, which is impractical and unfeasible as the number of necessary measurements scales exponentially with the number of qubits. Consequently, it is common practice to assume uncorrelated readout noise, allowing the multi-qubit noise matrix to be constructed as a tensor product of the single-qubit noise matrices as
@@
While making the tomographic step easier, even the multi-qubit noise matrix of Eq. (5) does not reduce the computational cost of implementing Eq. (4) and the Bayesian measurement error mitigation remains resource intensive and cannot be employed as it is. For all these reasons, we have developed a heuristic strategy that enables the use of the scheme in the multi-qubit case, as depicted in Fig. 1. The first key component involves reducing the number of populations $\rho_{\mathbf{i}}$ included in the iterative procedure. We achieve this by running the mitigation algorithm only on the subspace of multi-qubit state probabilities where noisy outcomes have been measured. In other words, we exclude from the Bayesian cycle any states for which no strings are obtained in the experiment, limiting it to the strings $\bm{i}$ for which $\rho_{\bm{i}}^{\mathrm{noisy}}\neq 0$. Subspace reduction is a common strategy for measurement error mitigation techniques which suffer from the same exponential scaling problem, as discussed in [36, 35].
@@
$\mathrm{Pr}(\bm{s}|\rho_{\bm{i}},\rho_{\bm{j}})=\Lambda_{\bm{s}\bm{i}}\rho_{\bm{i}}+\Lambda_{\bm{s}\bm{j}}\rho_{\bm{j}}+\sum_{\bm{k}\neq\bm{i},\bm{j}}\Lambda_{\bm{s}\bm{k}}R_{\bm{k}}.$ (6)
For a single measurement, Eq. (6) is still the conditional probability of measuring the multi-qubit string $\mathbf{s}$ when the correct probabilities for strings $\mathbf{i}$, $\mathbf{j}$, and the remaining $\mathbf{k}$ (with $\mathbf{i}\neq\mathbf{j}\neq\mathbf{k}$) are $\rho_{\bm{i}}$, $\rho_{\bm{j}}$, and $R_{\bm{k}}$ respectively. However, only the first two are treated as variables of the probability distribution, while the remaining ones are kept constant. Using this form for the conditional probability, we can apply the Bayesian iterative cycle in the same fashion trough
$\mathrm{Pr}(\rho_{\bm{i}},\rho_{\bm{j}}|\mathbf{s}^{(n)})\propto$ $\left[\Lambda_{\bm{s}\bm{i}}\rho_{\bm{i}}+\Lambda_{\bm{s}\bm{j}}\rho_{\bm{j}}+\sum_{\bm{k}\neq\bm{i},\bm{j}}\Lambda_{\bm{s}\bm{k}}R_{\bm{k}}\right]$ (7)
×$\times\mathrm{Pr}(\rho_{i},\rho_{j}|\mathbf{s}^{(n-1)}),$
and obtain a final probability distribution for the two variables. The posterior probability distribution $\mathrm{Pr}(\rho_{\bm{i}},\rho_{\bm{j}}|\mathbf{s}^{(\mathrm{N}_{\mathrm{shots}})})$ can then be used to estimate the correct values for the probabilities of measuring ${\bm{i}}$ and ${\bm{j}}$, thus obtaining some updated values for $R_{\bm{i}}$ and $R_{\bm{j}}$. After this step, the strategy is to repeat the procedure for a new pair of probabilities until all possible pair combinations have been updated once. This sequential heuristic binary optimization is then repeated until a chosen convergence criterion is achieved. A simple rule of thumb is to stop the algorithm repetitions if the change in the obtained populations does not change with respect to the previous iteration within a chosen accuracy. All the steps of the algorithm are outlined in Algorithm 2.
@@
$\mathrm{Pr}(\bm{Q}|\rho_{i},\rho_{j})=\Lambda_{\bm{i}}(\bm{Q})\rho_{\bm{i}}+\Lambda_{\bm{j}}(\bm{Q})\rho_{\bm{j}}+\sum_{\bm{k}\neq\bm{i},\bm{j}}\Lambda_{\bm{k}}(\bm{Q})R_{\bm{k}},$ (8)
@@
6:        Perform Bayesian cycle to obtain $\mathrm{Pr}(\rho_{i},\rho_{j})$

### v2

# Bayesian mitigation of measurement errors in multiqubit experiments

@@ Abstract
In Phys. Rev. A 108, L060402 (2023), we introduced a Bayesian measurement error mitigation algorithm, which leverages complete information from the readout signal, and validated the protocol on a quantum device with five superconducting qubits. Here, we present an improved algorithm’s implementation, tailored for multiqubit experiments on near-term superconducting qubit quantum devices. In particular, we provide a detailed algorithm workflow, from calibrating the detector response functions to the post-processing of measurement outcomes, offering a computationally efficient solution for the output size typical of current quantum computing devices. We show how the numerical representation of the noise function affects the performance of the error mitigation algorithm and test the convergence criteria. We benchmark our protocol on actual quantum computers with superconducting qubits, where the readout signal encodes the measurement information as unprocessed analog data before qubit state assignment. Finally, we compare the performance of our algorithm against other measurement error mitigation methods, such as iterative Bayesian unfolding and the Mthree method, and show how our method can be integrated on top of other readout error mitigation protocols.
@@ I Introduction
Despite rapid advancements in quantum computing technologies, practical applications of noisy intermediate-scale quantum (NISQ) processing units (QPUs) are still out of reach. While QPU manufacturers are constantly improving quality metrics [1, 2, 3], be it decoherence times, gate and readout fidelities, a lot of effort is being invested in developing software tools to shrink the gap between acceptable noise levels and utility-scale algorithms [4, 5]. Quantum error mitigation (QEM) embodies this goal, although a comprehensive and universally accepted definition for the field is lacking [6]. In general, QEM encompasses all those algorithmic schemes that reduce noise and errors in the expectation value of some figure of merit, such as the average of observables or quantum state probabilities. This is typically achieved by post-processing the output from an ensemble of circuit runs, either the same circuit or a family of properly designed ones [7]. This statistical approach starkly contrasts with quantum error correction, which needs to detect errors through single shot measurements. Among the most advanced techniques, zero-noise extrapolation [8] with probabilistic error amplification, probabilistic error cancellation [9], and tensor network error mitigation [10] have recently been found to be successful in the largest available QPUs [11, 12, 13, 14]. Some specialized techniques aim to target specific sources of errors, such as measurement error mitigation strategies. These strategies are designed to improve the readout step of quantum algorithms execution by using a model of the measurement error, which is a result of the faulty measurement process.
@@
Figure 1: (a) Workflow of standard readout error mitigation. (b) Workflow of Bayesian measurement error mitigation. (c) Workflow of Bayesian measurement error mitigation when integrating other readout error mitigation techniques. (d) Pictorial sketch of the heuristic algorithm to perform Bayesian measurement error mitigation in a multiqubit experiment. In the first step (I), we collect the noisy bitstrings counts and reduce the $2^{\mathrm{N_{q}}}$ outcome subspace to mitigate to the $M$ measured noisy counts for which we have at least one measurement. In step (II), we select two populations to use as variables while keeping all others constant, and apply Bayesian measurement error mitigation to obtain two new mitigated values. The Bayesian update is performed for all possible pairs. Step (II) is repeated until a specified convergence criterion or exit condition is met, resulting in the full set of mitigated values (III).
Most readout mitigation techniques are based on some form of detector or measurement tomography [15, 16, 17, 18] to obtain the assignment, noise, or confusion matrix. The readout error is then addressed in various ways. One straightforward strategy involves manipulating the outcome count statistics by multiplying the inverse of the confusion matrix by the measured probability vector of the quantum state [19]. However, this approach can produce nonphysical values, such as negative probabilities, an issue that can be taken into account by incorporating a convex optimizer into the mitigation pipeline [20]. Alternatively, some proposals have developed machine learning-based techniques to perform the assignment process [21, 22, 23, 24].
@@
In a recent work [25], we have introduced an efficient and accurate readout measurement scheme for single- and multi-qubit states based on Bayesian inference, inspired by successful application of Bayesian inference in quantum metrology and sensing [26, 27] or parameter estimation for quantum circuits and states [28, 29, 16, 30, 31, 32, 33, 34]. By bypassing the need to invert the noise matrix, our method leverages knowledge of the detector’s faulty response and employs Bayesian inference to estimate probability distributions for each qubit state. Moreover, such an approach is not limited to using binary outcomes from the detector, but it can exploit the detector unprocessed analog readout data before any binary assignment is made. The idea of using analog data has also found traction in quantum error correction, where it has been shown to lead to improved decoders and logical qubit error rates [35, 36]. Still, all measurement error mitigation schemes suffer some form of scalability issue when moving to many-qubit systems, as intuitively the noise matrix size scales exponentially with the number of qubits. Often, this issue is addressed by not using the full assignment matrix, or its inverse, and instead working in a subspace defined by the noisy outcomes [37, 38]. In this work, we aim to tackle thoroughly the scaling problem when aiming to mitigate measurement errors with our formerly introduced Bayesian readout and further show how the use of the readout analog data leads to increased performances. Beyond showing a solid proof of principle of the method, we make a full scaling analysis of the algorithm and provide an estimate of the classical computation time needed to post-process the measurement results obtained from currently available superconducting QPUs.
@@
The paper is organized as follows. In Section II, we present the theoretical foundations necessary to model measurement errors in quantum computers. We demonstrate how a Bayesian approach can be employed to infer the true outcomes from noisy measurements, assuming prior knowledge of the detector noise matrix. Additionally, we introduce a novel algorithm for implementing Bayesian measurement error mitigation in the context of multi-qubit measurements. The proposed method includes a detailed, step-by-step procedure suitable for current quantum hardware. We also describe how to adapt this framework to various types of measurement data, including binary outcomes (i.e., ”0”s and ”1”s) and unprocessed analog signals, such as IQ readout data for superconducting QPUs [39].
@@
In Section III, we benchmark our implementation of Bayesian measurement error mitigation by running quantum algorithms on actual quantum hardware. We demonstrate how the algorithm can be efficiently applied to improve readout accuracy in multi-qubit experiments and show how the performance of the mitigation scheme scales with both the number of qubits and the number of measurements for the chosen examples.
Section IV focuses on applying the Bayesian measurement error mitigation scheme to analog detector data. We show that utilizing the full IQ-readout signal—rather than the binary outcomes—can significantly improve the performance of the mitigation algorithm in multi-qubit scenarios. Additionally, we discuss the computational cost and convergence criteria of our implementation of the algorithm for multi-qubit Bayesian measurement error mitigation and provide, as an example, the computational times required for different state preparations. In Section V, we compare our Bayesian methods against two other measurement error mitigation schemes and show how it could be integrated with them. Throughout this work, we show the results obtained by executing the quantum circuits on three superconducting qubit devices: VTTQ20, VTTQ50, and IBM127 (Nazca). Finally, we present the conclusions and outline some possible future research directions and applications in Section VI.
@@ II Theoretical background
Let’s start with a mathematical description of a noisy measurement process for a single qubit. Let $s\in\{0,1\}$ be the measurement outcome observed in a single experiment. In the absence of measurement errors, the probability to register either one is fully determined by the quantum state of the system. Formally, a conventional single qubit measurement is modeled with a Positive-Operator Valued Measure (POVM) with two effects $\{M_{0}=|0\rangle\langle 0|,M_{1}=|1\rangle\langle 1|\}$. Consequently, the probability of observing a measurement outcome $s$, when the quantum system is in a state $\hat{\rho}$, can be calculated as $\mathrm{P}(s)=\operatorname{Tr}(\hat{\rho}M_{s})\equiv\rho_{s}$.
@@
However, in a real setup, the readout step is noisy and errors emerge in the process of collecting the signal from the detector and in the assignment of the correct qubit state. In practice, we can model the experimental noisy measurement through a set of noisy effects $M^{\mathrm{noisy}}_{i}=\sum_{j}\langle i|\Lambda|j\rangle M_{j}=\sum_{j}\Lambda_{ij}M_{j}$, where $\Lambda$ is the noise matrix and $\Lambda_{ij}$ is the probability of observing/assigning outcome $i$ when the true outcome is $j$. Therefore, after repeated measurements, we have direct access to the noisy probabilities
$\rho^{\mathrm{noisy}}_{s}=\Lambda_{sj}\rho_{j},$ (1)
@@
which are connected to the true qubit state probabilities through the noise matrix $\Lambda$. Thus, the goal of measurement error mitigation is to reconstruct the true qubit state probabilities after measuring $\rho^{\mathrm{noisy}}_{i}$, as in the workflow of Fig. 1 (a). Assuming that we have access to the noise matrix $\Lambda$, from Eq. (1), it is straightforward to attempt to invert the noise matrix and use it to reconstruct the correct qubit state probability vector as $\rho=\Lambda^{-1}\rho^{\mathrm{noise}}$. However, this simple measurement error mitigation strategy might result in a probability vector containing negative values due to statistical errors. This issue is often accounted for by employing a convex optimizer which project the corrected vector to a physical one [17, 20].
@@
To avoid these issues, we propose here a strategy to bypass the need to invert the noise matrix by noting how the measured outcome $\rho^{\mathrm{noisy}}_{s}$, as introduced in Eq. (1), can be read as the conditional probability of measuring $s$ when the correct qubit state probabilities are $\rho_{0}$ and $\rho_{1}$, i.e. $\rho^{\mathrm{noisy}}_{s}=\mathrm{P}(s|\rho_{0},\rho_{1})$.
@@
Thus, we can reconstruct a probability distribution for the correct probabilities $\rho_{0}$ and $\rho_{1}$ through Bayes’ theorem as $\mathrm{P}(\rho_{0},\rho_{1}|s)\propto\left[\Lambda_{s0}\rho_{0}+\Lambda_{s1}\rho_{1}\right]\mathrm{P}(\rho_{0},\rho_{1})$. It is convenient to follow an iterative approach and consider individually the $\mathrm{N}_{\mathrm{shots}}$ values collected after each measurement. In this way, we can update the conditional probability for the quantum state probabilities $\rho_{0}$ and $\rho_{1}$ shot after shot as
$\mathrm{P}(\rho_{0},\rho_{1}|s^{(n)})\propto\left[\Lambda_{s0}\rho_{0}+\Lambda_{s1}\rho_{1}\right]\mathrm{P}(\rho_{0},\rho_{1}|s^{(n-1)}),$ (2)
@@
where $s^{(n)}\in\{0,1\}$ is the bit registered in the n-th experiment, and $\mathrm{P}(\rho_{0},\rho_{1}|s^{(n-1)})$ is the prior probability distribution which corresponds to the posterior distribution calculated at the previous step. The idea is to update the prior probability distribution after each iteration step with the newly calculated posterior probability distribution. When assuming no prior knowledge about the system state, we start the iteration with a uniform prior probability distribution $\mathrm{Pr}(\rho_{0},\rho_{1}|s^{(0)})=\mathrm{const}$.
The end result of the Bayesian measurement error mitigation scheme is the probability distribution $\mathrm{P}(\rho_{0},\rho_{1}|s^{(\mathrm{N}_{\mathrm{shots}})})$ which represents the posterior probability distribution for the qubit state probabilities coherent with all the collected measurement data. Finally, this distribution can be used to estimate the ideal qubit state probabilities $\mathrm{P}{(0)}$ and $\mathrm{P}{(1)}$. The workflow of the mitigation algorithms follows Fig. 1 (b).
In our previous work [25], we proposed to use the first moment of the distribution $\mathrm{P}(\rho_{0},\rho_{1}|s^{(\mathrm{N}_{\mathrm{shots}})})$, namely $\mathrm{P}{(0)}=\int d\rho_{0}d\rho_{1}\;\rho_{0}\mathrm{P}(\rho_{0},\rho_{1}|s^{(\mathrm{N}_{\mathrm{shots}})})$ and similarly for $\mathrm{P}{(1)}$. Nevertheless, we found that results are highly improved by choosing the values $\rho_{0/1}$ for which the probability $\mathrm{P}(\rho_{0},\rho_{1}|s^{(\mathrm{N}_{\mathrm{shots}})})$ has its maximum , i.e. $(\mathrm{P}{(0)},\mathrm{P}{(1)})=\underset{\rho_{0},\rho_{1}}{\arg\max}\,\mathrm{P}(\rho_{0},\rho_{1}|s^{(\mathrm{N}_{\mathrm{shots}})})$.
@@
With these ingredients, we can extend the Bayesian measurement error mitigation scheme to the multi-qubit case with relative ease. For a system of $\mathrm{N}_{q}$ qubits, the measurement outcome obtained from a single experiment is a $\mathrm{N}_{q}$ digits string, $\mathbf{s}\in\{0,1\}^{\mathrm{N}_{q}}$. Equivalent to Eq. (1), the conditional probability of observing the multi-qubit string $\mathbf{s}$ is related to the correct outcome through a multi-qubit noise matrix as
$\mathrm{P}(\mathbf{s}|\bm{\rho}\equiv\rho_{0},...,\rho_{2^{N_{q}}-1})=\Lambda_{\mathbf{sj}}\rho_{\mathbf{j}}.$ (3)
@@
$\mathrm{P}(\bm{\rho}|\mathbf{s}^{(n)})\propto\Lambda_{\mathbf{s}^{(n)}\mathbf{j}}\rho_{\mathbf{j}}\mathrm{P}(\bm{\rho}|\mathbf{s}^{(n-1)}),$ (4)
@@
For example, for a system with $\mathrm{N_{q}}$} qubits, Eq. (4) requires handling a multi-parameter probability distribution $\mathrm{P}(\bm{\rho}|\mathbf{s}^{(n)})$ at each iteration step of the Bayesian update. This distribution involves $2^{\mathrm{N_{q}}}$ variables $\rho_{\mathbf{i}}$, each sampled within the interval $[0,1]$. If each variable is sampled with $n_{p}$ points, the total probability distribution would necessitate a mesh-grid with $n_{p}^{2^{\mathrm{N_{q}}}}$ points. Even though the normalization condition reduces the effective number of independent coordinates the computational cost remains high and out of reach.
Additional complications arise from the noise matrix resulting of the multi-qubit measurement process. Reconstructing the multi-qubit noise matrix in Eq. (4) ideally requires full detector tomography, but this quickly becomes infeasible as measurement requirements grow exponentially with qubit count. Consequently, it is common practice to assume uncorrelated readout noise, allowing the multi-qubit noise matrix to be constructed as a tensor product of the single-qubit noise matrices as
@@
While making the tomographic step easier, even the multi-qubit noise matrix of Eq. (5) does not reduce the computational cost of implementing Eq. (4) and the Bayesian measurement error mitigation remains resource intensive and cannot be employed as it is. For all these reasons, we have developed a heuristic strategy that enables the use of the scheme in the multi-qubit case, as depicted in Fig. 1 (d). The first key component involves reducing the number of populations $\rho_{\mathbf{i}}$ included in the iterative procedure. We achieve this by running the mitigation algorithm only on the subspace of multi-qubit state probabilities where noisy outcomes have been measured. In other words, we exclude from the Bayesian cycle any states for which no strings are obtained in the experiment, limiting it to the strings $\bm{i}$ for which $\rho_{\bm{i}}^{\mathrm{noisy}}\neq 0$. Subspace reduction is a common strategy for measurement error mitigation techniques which suffer from the same exponential scaling problem, as discussed in [40, 37].
@@
$\mathrm{P}(\bm{s}|\rho_{\bm{i}},\rho_{\bm{j}})=\Lambda_{\bm{s}\bm{i}}\rho_{\bm{i}}+\Lambda_{\bm{s}\bm{j}}\rho_{\bm{j}}+\sum_{\bm{k}\neq\bm{i},\bm{j}}\Lambda_{\bm{s}\bm{k}}R_{\bm{k}}.$ (6)
For a single measurement, Eq. (6) is still the conditional probability of measuring the multi-qubit string $\mathbf{s}$ when the correct probabilities for strings $\mathbf{i}$, $\mathbf{j}$, and the remaining $\mathbf{k}$ (with $\mathbf{i}\neq\mathbf{j}\neq\mathbf{k}$) are $\rho_{\bm{i}}$, $\rho_{\bm{j}}$, and $R_{\bm{k}}$ respectively. However, only the first two are treated as variables of the probability distribution, while the remaining ones are kept constant. Using this form for the conditional probability, we can apply the Bayesian iterative cycle in the same fashion through
$\mathrm{P}(\rho_{\bm{i}},\rho_{\bm{j}}|\mathbf{s}^{(n)})\propto$ $\left[\Lambda_{\bm{s}\bm{i}}\rho_{\bm{i}}+\Lambda_{\bm{s}\bm{j}}\rho_{\bm{j}}+\sum_{\bm{k}\neq\bm{i},\bm{j}}\Lambda_{\bm{s}\bm{k}}R_{\bm{k}}\right]\$(7)
$\times\mathrm{P}(\rho_{i},\rho_{j}|\mathbf{s}^{(n-1)}),$
and obtain a final probability distribution for the two variables. The posterior probability distribution $\mathrm{P}(\rho_{\bm{i}},\rho_{\bm{j}}|\mathbf{s}^{(\mathrm{N}_{\mathrm{shots}})})$ can then be used to estimate the correct values for the probabilities of measuring $\bm{i}$} and $\bm{j}$}, thus obtaining some updated values for $R_{\bm{i}}$ and $R_{\bm{j}}$. Furthermore, throughout our work, we sample the probability distributions for $\rho_{\bm{i}}$ and $\rho_{\bm{j}}$ on a grid with $n_{p}=R_{\bm{i}}+R_{\bm{j}}$ points. After this step, the strategy is to repeat the procedure for a new pair of probabilities until all possible pair combinations have been updated once. It is possible to further reduce the numerical complexity by considering pairs whose bit strings are within Hamming distance $d$, e.g. $H_{d}(i,j)\leq d$. This sequential heuristic binary optimization is then repeated until a chosen convergence criterion is achieved. A simple rule of thumb is to stop the algorithm repetitions if the change in the obtained populations does not change with respect to the previous iteration within a chosen accuracy. All the steps of the algorithm are outlined in Algorithm II.
@@
$\mathrm{P}(\bm{Q}|\rho_{i},\rho_{j})=\Lambda_{\bm{i}}(\bm{Q})\rho_{\bm{i}}+\Lambda_{\bm{j}}(\bm{Q})\rho_{\bm{j}}+\sum_{\bm{k}\neq\bm{i},\bm{j}}\Lambda_{\bm{k}}(\bm{Q})R_{\bm{k}},$ (8)
@@
6:   Perform Bayesian cycle to obtain $\mathrm{P}(\rho_{i},\rho_{j})$